OpenStudy (anonymous):

How do you find the percentage of something? Usually I just put it into a calculator like: 16 percent of 64 - I would just put 16 * 64 = 1024 so I would assume that it is 10.24 But now I have to actually show my work for it and I don't understand what they mean by something like: what percent of 42 is 70? I have no clue what I am doing now!

hartnn (hartnn):

x% = x/100

hartnn (hartnn):

let x% of 42 =70 so, x%*42 =70 x*42/100 = 70

hartnn (hartnn):

solve for x now.

Directrix (directrix):

@haileemackk >16 percent of 64 - I would just put >16 * 64 = 1024 so I would assume that it is 10.24 The above process is not correct. Maybe you did 16/64 and got 25 and assumed it was 25%. ------------ what percent of 42 is 70? Percent means hundredths. That means 42 is to 70 as x is to 100. 42/70 = x/100 x = 60 Looking back at the proportion, x is over 100. So, 60/100 = 60% is the percentage.

hartnn (hartnn):

ummm....60% of 70 is 42 but 60% of 42 is not 70(i think this is asked)

hartnn (hartnn):

16 percent of 64 = 16% *64 = 16/100 *64 = 10.24